# ColorBench: Can VLMs See and Understand the Colorful World? A Comprehensive Benchmark for Color Perception, Reasoning, and Robustness

### v2

@@ 1 Introduction
However, existing benchmarks for VLMs mainly focus on tasks that may not heavily depend on color understanding or require color-centric reasoning, thereby overlooking nuanced color-related factors [11, 12]. Hence, there is a lack of benchmarks that systematically assess how well VLMs understand color when it serves as the main or distinguishing feature of a scene and key information to a task. Moreover, robustness to variations in color, such as recoloring and shifting hues, has also been largely neglected in the LLM era [13, 14, 15]. Consequently, it remains unclear whether VLMs can perceive and reason about color with human-like proficiency and to what extent their performance deteriorates under significant color perturbations. This shortfall underscores the need for a dedicated benchmark that comprehensively probes various facets of color comprehension in VLMs.
@@
By focusing on these facets, ColorBench offers a granular view of VLMs’ capabilities in color understanding, aiming to illuminate both their strengths and shortcomings. We evaluate $32$ widely used VLMs in our benchmark, ranging from open-source to proprietary models, from relatively small models ($0.5$B) to larger models ($78$B), and obtain some unrevealed observations.
@@ 2 ColorBench Construction
Figure 2: Statistics of 3 categories and 11 tasks in ColorBench.
@@ 2.1.2 Color Reasoning
Color Proportion tests the model’s capability to estimate the relative area occupied by a specific color. Questions in this task require both color perception and proportion calculation capabilities. Color Comparison requires the model to be able to distinguish among multiple colors in the image, assessing its sensitivity to hue, saturation, and brightness differences in visual input. Color Counting focuses on identifying the number of unique colors in the image, evaluating the model’s perception and differentiation of distinct color variations, and counting ability. Object Counting extends this challenge by requiring the model to count objects that match a specific color pattern. This task requires an integration ability of object recognition and color perception. Color Illusion questions query VLMs to compare colors in potential illusionary environments. This task evaluates the model’s ability to account for color-induced optical illusions. Color Mimicry challenges the model to detect objects camouflaged within their surroundings, where color serves as a misleading factor, requiring advanced pattern recognition and contextual reasoning. These two tasks both assess the model’s ability to make correct predictions under the misleading of color-related information in visual input. Color Blindness, inspired by Ishihara tests, assesses the model’s ability to recognize numbers or text embedded in color patterns, testing its understanding of shape-color relationships. These $7$ tasks comprehensively assess the model’s capacity for logical reasoning, spatial awareness, and adaptive interpretation of color-based visual cues.
@@ 2.1.3 Color Robustness
Figure 3: Generation Pipeline for Color Robustness. For each seed image, we apply $3$ recoloring strategies (Entire Image, Target Segment, Largest Segment) to generate edited images. For each strategy, we change the color of the recoloring region via shifting the Hue values by $90$°, $180$°, or $270$°in HSV color space.
@@ 2.2 Data Curation
For most of the tasks in the category of Color Perception and Color Reasoning, we rely on human experts to manually collect images from multiple online benchmarks and websites. For the Color Proportion task, to ensure the correctness of the ground truth, an extra color extraction tool is firstly utilized to obtain the color histogram of the image. Questions and options are then manually designed based on these color statistics. For tasks including Color Extraction, Color Blindness, and Color Illusion, testing images are generated by corresponding code programs to ensure the controllability of the questions and answers. The detailed data sources are shown in Appendix A.
After the initial data is collected, additional filtering processes are conducted in a human-machine interactive process. We first conduct inference on a variety of VLMs and discard low-quality samples based on GPT-4o prediction result and human evaluation. For synthesized data, similar processes are conducted but with additional code (for generation) and image assessment. The above process is conducted in three rounds before the final benchmark instances are settled. This refinement process ensures ColorBench a rigorous and informative benchmark for assessing color-related understanding.
For Color Robustness, we create evaluation instances by modifying images or specific regions through color changes. We define $3$ recoloring strategies to determine the recoloring region: i) Entire Image, where the whole image is recolored; ii) Target Segment, where only the segment relevant to the question is altered; and iii) Largest Segment, where the largest region unrelated to the question is modified. Further details can be found in Appendix B. While generating color variants, we derive seed images from CV-Bench [16], a publicly available benchmark. For each seed image, as shown in Figure 3, we first employ a Grounded Segmentation Model (GAM) [17] to extract segments and their corresponding labels. We then apply the predefined recoloring strategies to determine the editing region and perform recoloring by shifting the Hue value in the HSV color space at three levels to cover entire color wheel: ($90$°, $180$°, and $270$°). This process produces $9$ variations per seed image, covering different strategies and degrees of color change to enable a comprehensive robustness assessment. To ensure interpretability, human experts filter out unnatural or negligible modifications, resulting in a final selection of $493$ seed images for robustness evaluation.
@@ 2.3 Evaluation Metrics
Instance-level Robustness. Let the recolored images be $I_{1},\cdots,I_{n}$ and the generation output of model for image $I_{i}$ and question $q$ is $\mathcal{M}(I_{i},q)$. Define $c(\mathcal{M}(I_{i},q))$ as the model correctness: $c(\mathcal{M}(I_{i},q))=1$ if model result $\mathcal{M}(I_{i},q)$ is correct, otherwise 00. The instance-level robustness metric $R(I_{s},q)$ for a seed image $I_{s}$ and question $q$ is defined as:
@@ 3.1 Main Results
Table 1 presents the performances of a wide range of VLMs, along with human evaluation results on our ColorBench. Human participants achieve the highest performance on all evaluated tasks across all models. Among the models, overall accuracy generally increases with model size, with larger models tend to outperform smaller models, and the two proprietary models, GPT-4o and Gemini-2-flash, perform the best [^1] To examine the upper limits of VLM capabilities and benchmark against human-level performance, we also assess performance GPT-o3 on perception and reasoning tasks. The result is shown in Appendix G..
@@
Color Robustness. In Color Robustness (C’Robust), a higher value represents better robustness towards color alteration. The only $4$ models that exceed $80\%$ are LLaVA-OV-72B, InternVL2.5-26B, InternVL2.5-38B, and InternVL2.5-78B, which utilize relatively larger vision encoders, InternViT-6B, compared with others (mostly only 300-400M). In the meantime, GPT-4o has a really low robustness ($46.2\%$) to colors, indicating its vulnerable sensitivity to color changes, while Gemini-2 shows promising robustness ($70.7\%$) towards colors. Moreover, another surprising observation is that even though only the colors are changed and all the original queries are kept, utilizing more reasoning steps can consistently improve robustness for GPT-4o ($+23.7\%$) and Gemini-2 ($+2.9\%$).
@@ 3.2 Further Findings
Since color-related tasks often involve abstract reasoning, language comprehension, and contextual interpretation, it is essential to assess not just the vision encoder but also part of the language model, which plays a critical role in processing and understanding such tasks. To quantitatively analyze the correlation between VLM performances on color understanding tasks and their sizes, Spearman’s rank correlation is calculated between VLM performances and (i) overall model sizes (L $+$ V), (ii) language model sizes (L), and (iii) vision encoder sizes (V). The correlation values and p signs are presented in Table 2; a star is notated when the p-value of the correlation is lower than $0.05$. It is observed that between the performances and language model sizes, most of the tasks have a correlation greater than $0.5$ and a p-value smaller than $0.05$, except for Color Illusion and Color Blindness due to their special characteristics. Since the correlation between overall model sizes (L $+$ V) and P&R Overall ($0.7619$), and Robustness ($0.7390$), we conclude that the color understanding, including Color Perception, Color Reasoning, and Color Robustness, still follows the scaling law of model sizes. Figure 4 presents the correlations between performances and model sizes in each model family. This visualization clearly shows the correlation between performances and model sizes; a larger model leads to higher performance within each model family.
(a) VLM Size & Model Performance
(b) Vision Size & Model Performance
Figure 5: The scatter plots related to performances and model sizes. Each plot illustrates the relationship between the log-scaled size of parameter numbers and the performance across all models. (a) Number of parameters for the entire VLM; (b) Number of parameters for the vision encoder. There is a significant correlation between performance and VLM sizes. However, since the choices of vision encoder size are concentrated and limited, it is hard to conclude the relationship between performance and vision sizes.
However, between the performances and vision encoder sizes, most of the tasks either have a correlation lower than $0.5$ or a p-value greater than $0.05$, which is not sufficient to conclude with evident positive correlation. Figure 5(b) visualizes the correlations between model performances and vision encoder’s sizes for all models. Despite these findings, we try to avoid conveying the message that there is no positive correlation between performances and vision encoder sizes. We think it is because of the negligence of the current community to focus on the scaling laws of vision encoders. The vision encoders used in the current mainstream VLMs are constrained in a very small set: (i) most of the VLMs only use one type of vision encoders for the whole family, except for the InternVL2 and InternVL2.5 series; (ii) most of VLMs use the vision encoder with the size of $300$ $400$M. These challenges make it hard to evaluate the scaling laws of vision encoders. Further visualizations are presented in Appendix H.2.
@@
The absolute performances of different VLMs are relatively low and lag behind that of humans. Moreover, the gaps between different models (open-source vs. proprietary, small vs. large) are not large, indicating the challenges of ColorBench and negligence of color understanding in existing VLMs.
@@
As shown in Table 3, we separate all the VLMs into several groups based on their sizes and present the best accuracy and the model name within each group. We can see that even the powerful proprietary models, GPT-4o and Gemini-2, can only reach an overall color perception and reasoning (P & R Overall) accuracy of $53.9\%$, only $+2.0\%$ better than the best open-sourced model. Task-level results in Table 1 further reveal that these advanced proprietary models still exhibit substantial performance gaps compared to humans across most tasks. Moreover, the best model from group 1 has the accuracy of $41.5\%$ (Cambrian-3B), which is only $10.4\%$ lower than the best open-sourced model. As for the robustness, the powerful proprietary models even show weaker robustness than the 7B model. Considering the lack of existing benchmarks specifically evaluating VLMs’ color understanding capabilities, we conclude that this area is long-neglected by the community, and the open-sourced community is still on the same page with the proprietary model providers.
@@
The impact of using CoT prompting is shown in Table 4, in which we can see CoT improves the average P&R Overall accuracy across both models by $+4.35\%$, indicating that reasoning benefits these color-related tasks. Within the category of Color Perception, the improvements from CoT on Color Recognition and Object Recognition are quite limited as these tasks heavily rely on the vision encoder. Figure 51 and 52 in Appendix I illustrate that adding reasoning steps does not take effect since the initial visual perception and color identification are incorrect in the slow thinking process. However, to our surprise, we find that the Color Extraction task benefits extremely from more reasoning steps, although it seems only related to the vision encoder. After a thorough investigation, we observe that most of the current VLMs are not capable of directly extracting color values, so they need to use more reasoning steps to reach reasonable answers.
@@
In order to examine whether VLMs really leverage color clues to handle tasks in ColorBench, experiments are conducted by converting all the original colorful images in the Color Perception and Reasoning categories into gray-scale ones, without changing the questions. Under this circumstance, the accuracies are expected to decrease dramatically as all our questions are related to colors. For quantitative analysis, we calculate the accuracy changing ratio as $(Acc_{ori}-Acc_{gray})/Acc_{ori}$ for each VLM on each task. This value directly represents how the original accuracy changes with a gray-scale transformation. The positive value represents that the VLM has a higher accuracy on the original colored images, indicating that it needs color clues to solve the task. Higher positive values represent higher significance of the color clues. On the contrary, if the value is negative, it means that the VLM can reach a better accuracy after the gray-scale transformation, indicating that it does not need color clues for the task, and colors might even mislead VLM’s judgment. Lower negative values represent severe harm the color can have on the task.
The accuracy changing ratio distribution across all VLMs and tasks are presented in Figure 6 as the violin plot. As shown in the figure, for most of the tasks, the ratios of VLMs are above 00, indicating that VLMs indeed leverage color clues to correctly solve the tasks; removing the color directly harms the original accuracies dramatically. However, when it comes to Color Illusion and Color Mimicry, the majority of the changing ratios are beneath 00, which means that VLMs can get better accuracies when all the color information is removed. This phenomenon is reasonable as the colors on both of these two tasks are more likely serving as the misleading factors. In the meantime, for the Color Counting and Color Blindness tasks, almost half the accuracies increase and half decrease, indicating that the color clues might not be so significant in this task, thus, some of the models can find other ways to get the answer. We also investigate the correlation between accuracy changing ratios and model sizes, while no significant correlation can be concluded.
4 Related Works
4.1 VLM Benchmarks

### v3

@@ 1 Introduction
However, existing benchmarks for VLMs mainly focus on tasks that may not heavily depend on color understanding or require color-centric reasoning, thereby overlooking nuanced color-related factors [25, 29]. Hence, there is a lack of benchmarks that systematically assess how well VLMs understand color when it serves as the main or distinguishing feature of a scene and key information to a task. Moreover, robustness to variations in color, such as recoloring and shifting hues, has also been largely neglected in the LLM era [6, 8, 20]. Consequently, it remains unclear whether VLMs can perceive and reason about color with human-like proficiency and to what extent their performance deteriorates under significant color perturbations. This shortfall underscores the need for a dedicated benchmark that comprehensively probes various facets of color comprehension in VLMs. A detailed discussion of related works is provided in Appendix A.
@@
By focusing on these facets, ColorBench offers a granular view of VLMs’ capabilities in color understanding, aiming to illuminate both their strengths and shortcomings. We evaluate $32$ widely used VLMs in our benchmark, ranging from open-source to proprietary models, from relatively small models (0.50.5B) to larger models (7878B), and obtain some unrevealed observations.
@@ 2 ColorBench Construction
Figure 2: Statistics of 3 categories and $1$ tasks in ColorBench.
@@ 2.1.2 Color Reasoning
Color Proportion tests the model’s capability to estimate the relative area occupied by a specific color. Questions in this task require both color perception and proportion calculation capabilities. Color Comparison requires the model to be able to distinguish among multiple colors in the image, assessing its sensitivity to hue, saturation, and brightness differences in visual input. Color Counting focuses on identifying the number of unique colors in the image, evaluating the model’s perception and differentiation of distinct color variations, and counting ability. Object Counting extends this challenge by requiring the model to count objects that match a specific color pattern. This task requires an integration of object recognition and color perception. Color Illusion questions query VLMs to compare colors in potential illusionary environments. This task evaluates the model’s ability to account for color-induced optical illusions. Color Mimicry challenges the model to detect objects camouflaged within their surroundings, where color serves as a misleading factor, requiring advanced pattern recognition and contextual reasoning. These two tasks both assess the model’s ability to make correct predictions under the misleading of color-related information in visual input. Color Blindness, inspired by Ishihara tests, assesses the model’s ability to recognize numbers or text embedded in color patterns, testing its understanding of shape-color relationships. These $7$ tasks comprehensively assess the model’s capacity for logical reasoning, spatial awareness, and adaptive interpretation of color-based visual cues.
@@ 2.1.3 Color Robustness
Figure 3: Generation Pipeline for Color Robustness. For each seed image, we apply 33 recoloring strategies (Entire Image, Target Segment, Largest Segment) to generate edited images. For each strategy, we change the color of the recoloring region via shifting the Hue values by 9090°, 180180°, or 270270°in HSV color space.
@@ 2.2 Data Curation
For most of the tasks in the category of Color Perception and Color Reasoning, we rely on human experts to manually collect images from multiple online benchmarks and websites. For the Color Proportion task, to ensure the correctness of the ground truth, an extra color extraction tool is firstly utilized to obtain the color histogram of the image. Questions and options are then manually designed based on these color statistics. For tasks including Color Extraction, Color Blindness, and Color Illusion, testing images are generated by corresponding code programs to ensure the controllability of the questions and answers. The detailed data sources are shown in Appendix B.
After the initial data is collected, additional filtering processes are conducted in a human-machine interactive process. We first conduct inference on a variety of VLMs and discard low-quality samples based on the GPT-4o prediction result and human evaluation. For synthesized data, similar processes are conducted, but with additional code (for generation) and image assessment. The above process is conducted in three rounds before the final benchmark instances are settled. This refinement process ensures ColorBench a rigorous and informative benchmark for assessing color-related understanding.
For Color Robustness, we create evaluation instances by modifying images or specific regions through color changes. We define $3$ recoloring strategies to determine the recoloring region: i) Entire Image, where the whole image is recolored; ii) Target Segment, where only the segment relevant to the question is altered; and iii) Largest Segment, where the largest region unrelated to the question is modified. Further details can be found in Appendix C. While generating color variants, we derive seed images from CV-Bench [42], a publicly available benchmark. For each seed image, as shown in Figure 3, we first employ a Grounded Segmentation Model (GAM) [38] to extract segments and their corresponding labels. We then apply the predefined recoloring strategies to determine the editing region and perform recoloring by shifting the Hue value in the HSV color space at three levels to cover entire color wheel: (9090°, 180180°, and 270270°). This process produces $9$ variations per seed image, covering different strategies and degrees of color change to enable a comprehensive robustness assessment. To ensure interpretability, human experts filter out unnatural or negligible modifications, resulting in a final selection of $493$ seed images for robustness evaluation.
@@ 2.3 Evaluation Metrics
Instance-level Robustness. Let the recolored images be $I_{1},\cdots,I_{n}$ and the generation output of model for image $I_{i}$ and question $q$ is $\mathcal{M}(I_{i},q)$. Define $c(\mathcal{M}(I_{i},q))$ as the model correctness: $c(\mathcal{M}(I_{i},q))=1$ if model result $\mathcal{M}(I_{i},q)$ is correct, otherwise 0. The instance-level robustness metric $R(I_{s},q)$ for a seed image $I_{s}$ and question $q$ is defined as:
@@ 3.1 Main Results
Table 1 presents the performances of a wide range of VLMs, along with human evaluation results on our ColorBench. Human participants achieve the highest performance on all evaluated tasks across all models. Among the models, overall accuracy generally increases with model size, with larger models tend to outperform smaller models, and the two proprietary models, GPT-4o and Gemini-2-flash, perform the best [^1] To examine the upper limits of VLM capabilities and benchmark against human-level performance, we also assess performance GPT-o3 on perception and reasoning tasks. The result is shown in Appendix H..
@@
Color Robustness. In Color Robustness (C’Robust), a higher value represents better robustness towards color alteration. The only $4$ models that exceed $80\%$ are LLaVA-OV-72B, InternVL2.5-26B, InternVL2.5-38B, and InternVL2.5-78B, which utilize relatively larger vision encoders, InternViT-6B, compared with others (mostly only 300-400M). In the meantime, GPT-4o has a really low robustness ($46.2\%$) to colors, indicating its vulnerable sensitivity to color changes, while Gemini-2 shows promising robustness ($70.7\%$) towards colors. Moreover, another surprising observation is that even though only the colors are changed and all the original queries are kept, utilizing more reasoning steps can consistently improve robustness for GPT-4o ($23.7\%$) and Gemini-2 ($2.9\%$).
@@ 3.2 Further Findings
The scaling law still holds for color understanding, but is much weaker and mainly depends on the language model parts. The correlation between the performance and the vision encoder’s size is not significant due to the limited choices in current VLMs.
@@
Since color-related tasks often involve abstract reasoning, language comprehension, and contextual interpretation, it is essential to assess not just the vision encoder but also part of the language model, which plays a critical role in processing and understanding such tasks. To quantitatively analyze the correlation between VLM performances on color understanding tasks and their sizes, Spearman’s rank correlation is calculated between VLM performances and (i) overall model sizes (L ++ V), (ii) language model sizes (L), and (iii) vision encoder sizes (V). The correlation values and p-signs are presented in Table 2; a star is notated when the p-value of the correlation is lower than $0.05$. It is observed that between the performances and language model sizes, most of the tasks have a correlation greater than $0.5$ and a p-value smaller than $0.05$, except for Color Illusion and Color Blindness due to their special characteristics. Since the correlation between overall model sizes (L ++ V) and P&R Overall ($0.7619$), and Robustness ($0.7390$), we conclude that the color understanding, including Color Perception, Color Reasoning, and Color Robustness, still follows the scaling law of model sizes. Figure 4 presents the correlations between performances and model sizes in each model family. This visualization clearly shows the correlation between performances and model sizes; a larger model leads to higher performance within each model family.
However, between the performances and vision encoder sizes, most of the tasks either have a correlation lower than $0.5$ or a p-value greater than $0.05$, which is not sufficient to conclude with the evident positive correlation. Despite these findings, we try to avoid conveying the message that there is no positive correlation between performances and vision encoder sizes. We think it is because of the negligence of the current community to focus on the scaling laws of vision encoders. The vision encoders used in the current mainstream VLMs are constrained in a very small set: (i) most of the VLMs only use one type of vision encoders for the whole family, except for the InternVL2 and InternVL2.5 series; (ii) most of the VLMs use the vision encoder with the size of $300$ - 400400M. These challenges make it hard to evaluate the scaling laws of vision encoders. Further visualizations are presented in Appendix L.2.
@@
The absolute performances of different VLMs are relatively low and lag behind those of humans. Moreover, the gaps between different models (open-source vs. proprietary, small vs. large) are not large, indicating the challenges of ColorBench and the negligence of color understanding in existing VLMs.
@@
As shown in Table 3, we separate all the VLMs into several groups based on their sizes and present the best accuracy and the model name within each group. We can see that even the powerful proprietary models, GPT-4o and Gemini-2, can only reach an overall color perception and reasoning (P & R Overall) accuracy of $53.9\%$, only $2.0\%$ better than the best open-sourced model. Task-level results in Table 1 further reveal that these advanced proprietary models still exhibit substantial performance gaps compared to humans across most tasks. Moreover, the best model from group 1 has the accuracy of $41.5\%$ (Cambrian-3B), which is only $10.4\%$ lower than the best open-sourced model. As for the robustness, the powerful proprietary models even show weaker robustness than the 7B model. Considering the lack of existing benchmarks specifically evaluating VLMs’ color understanding capabilities, we conclude that this area is long-neglected by the community, and the open-sourced community is still on the same page with the proprietary model providers.
@@
The impact of using CoT prompting is shown in Table 4, in which we can see CoT improves the average P&R Overall accuracy across both models by $4.35\%$, indicating that reasoning benefits these color-related tasks. Within the category of Color Perception, the improvements from CoT on Color Recognition and Object Recognition are quite limited as these tasks heavily rely on the vision encoder. Figure 59 and 60 in Appendix M illustrate that adding reasoning steps does not take effect since the initial visual perception and color identification are incorrect in the slow thinking process. However, to our surprise, we find that the Color Extraction task benefits extremely from more reasoning steps, although it seems only related to the vision encoder. After a thorough investigation, we observe that most of the current VLMs are not capable of directly extracting color values, so they need to use more reasoning steps to reach reasonable answers.
@@
In order to examine whether VLMs really leverage color clues to handle tasks in ColorBench, experiments are conducted by converting all the original colorful images in the Color Perception and Reasoning categories into gray-scale ones, without changing the questions. Under this circumstance, the accuracies are expected to decrease dramatically as all our questions are related to colors. For quantitative analysis, we calculate the accuracy changing ratio as $(Acc_{ori}-Acc_{gray})/Acc_{ori}$ for each VLM on each task. This value directly represents how the original accuracy changes with a gray-scale transformation. The positive value represents that the VLM has a higher accuracy on the original colored images, indicating that it needs color clues to solve the task. Higher positive values represent higher significance of the color clues. On the contrary, if the value is negative, it means that the VLM can reach a better accuracy after the gray-scale transformation, indicating that it does not need color clues for the task, and colors might even mislead VLM’s judgment. Lower negative values represent the severe harm the color can have on the task.
The accuracy changing ratio distributions across all VLMs and tasks are presented in Figure 5 as the violin plot. As shown in the figure, for most of the tasks, the ratios of VLMs are above 0, indicating that VLMs indeed leverage color clues to correctly solve the tasks; removing the color directly harms the original accuracies dramatically. However, when it comes to Color Illusion and Color Mimicry, the majority of the changing ratios are below 0, which means that VLMs can get better accuracies when all the color information is removed. This phenomenon is reasonable as the colors on both of these two tasks are more likely serving as the misleading factors. In the meantime, for the Color Counting and Color Blindness tasks, almost half the accuracies increase and half decrease, indicating that the color clues might not be so significant in this task, thus, some of the models can find other ways to get the answer. We also investigate the correlation between accuracy changing ratios and model sizes, while no significant correlation can be concluded.
4 Conclusion, Limitation, and Future Works
In this paper, we introduce ColorBench, the first benchmark designed to comprehensively evaluate the color understanding capabilities of VLMs, including Perception, Reasoning, and Robustness. After evaluating $32$ widely used VLMs on our benchmark, several undiscovered observations are revealed by us. These observations emphasize the need for more sophisticated model architectures that integrate deeper color reasoning capabilities. To ensure high-quality and reliable annotations, ColorBench relies on manual data collection, annotation, and assessment across most domains. While this guarantees consistency, it inevitably limits dataset scale, style diversity, and category coverage. As future work, we aim to develop a trustworthy automated data collection pipeline and expand ColorBench to larger-scale, more diverse tasks involving complex interplays of color with texture, shape, and spatial relationships. Furthermore, investigating the impact of different visual encoders and language models could further elucidate the pathways through which VLMs process color information.
Appendix A Related Works
A.1 VLM Benchmarks